# Quasiparticle gap renormalization driven by internal and external screening in WS2 device

### v1

@@ Abstract
Figure 1: (a) Schematic of microARPES experiment on a gated SL WS2 device. Gate and drain electrodes are labelled as G and D. (b) Optical image of the device. (c) Energy- and momentum-integrated map of ARPES intensity from the device region. The red triangular outline and the blue square outline in (b)-(c) demarcate the SL WS2 and graphene flakes, respectively. (d) ARPES spectra of the graphene Dirac cone at the given gate voltages ($V_{G}$) obtained along the cut through $\mathrm{K}_{\mathrm{Gr}}$ marked on the BZs of graphene (black hexagon) and SL WS2 (red hexagon) in the inset. The measurement was performed in the spot marked by a blue star in (c). (e) Graphene carrier density $n_{\mathrm{Gr}}$ as a function of $V_{\mathrm{G}}$ obtained from ARPES spectra of the Dirac cone as shown in (d) [34]. Red full lines represent linear fits with extrapolation shown as red dashed lines. The saturation level for electron-doping of the graphene is demarcated by a blue dashed line. (f)-(g) ARPES spectra of SL WS2 at the stated values of $V_{G}$ measured along $\mathrm{\Gamma}-\mathrm{K}_{\mathrm{W}}$. The ARPES cut direction is shown on the BZs in the inset. The energies of the VBM and CBM obtained from EDC fits are marked by red and blue ticks for spectra obtained in the two regions of the device demarcated by (f) red and (g) blue stars in (b)-(c) [34].
@@
Figure 2: (a)-(b) Close-up of ARPES spectra of the CBM region in (a) vacuum exposed and (b) graphene-encapsulated SL WS2 with estimates of electron density $n_{\mathrm{W}}$. (c) EDCs extracted along the dashed vertical line at $\mathrm{K}_{\mathrm{W}}$ in (a)-(b). Fits to a function composed of a single Lorentzian peak on a linear background multiplied by the Fermi-Dirac distribution are overlaid. The blue and red ticks demarcate the fitted peak positions. Double-headed arrows demarcate the full width at half maximum of the fitted Lorentzian peaks. Values are stated in units of eV and have error bars of ±$\pm 0.02$ eV.
@@
Turning to the vacuum exposed section of SL WS2 in the device, we present the $E(k)$ band dispersion along $\mathrm{\Gamma}-\mathrm{K}_{\mathrm{W}}$ in Fig. 1(f). Note that the Fermi level is always referenced to the Fermi energy of the graphene top contact. We present spectra at $V_{\mathrm{G}}=0$ V and at the electron-doping saturation level of $V_{\mathrm{G}}=8$ V determined on graphene. At $V_{\mathrm{G}}=0$ V, we observe the valence band maximum (VBM) to be located at $\mathrm{K}_{\mathrm{W}}$ at an energy of -1.48 eV and with a spin-orbit splitting of 0.42 eV, consistent with the expected band structure of SL WS2 [36]. At $V_{\mathrm{G}}=8$ V, the VBM is shifted by 0.61 eV to lower energies, the CBM has shifted below $E_{\mathrm{F}}$ and a substantial energy broadening of the bands has occurred. We find that the linewidth of energy distribution curves (EDCs) through the VBM nearly doubles compared to the situation at $V_{\mathrm{G}}=0$ V [34]. Such broadening primarily results from photocurrents in the SL WS2, which cause inhomogeneous doping and energy-shifting of the bands on the 4 $\mu$m length scale of the light spot [27, 30, 31].
@@
We do not encounter this problem in the graphene-encapsulated section of the SL WS2 device, as we do not observe any energy broadening under the graphene [34]. ARPES spectra are presented in Fig. 1(g). The reduced intensity and higher background compared with the spectra of bare SL WS2 in Fig. 1(f) are caused by inelastic photoelectron scattering in the overlying graphene. At $V_{\mathrm{G}}=0$ V, the VBM is located at -1.40 eV and thus shifted further towards $E_{\mathrm{F}}$ than in the vacuum exposed SL WS2 section of the device. This shift is due to the enforced band alignment upon interfacing SL WS2 with graphene. Tuning to $V_{\mathrm{G}}=12$ V leads to a downwards VBM shift of 0.80 eV and a faint CBM state at $E_{\mathrm{F}}$. We do not observe additional energy and momentum broadening of the SL WS2 bands in this part of the device when we increase $V_{\mathrm{G}}$ to the maximum possible value of 12 V before gate leakage through hBN sets in [34].
@@
Detailed ARPES spectra of the SL WS2 CBM region are presented in Fig. 2 for the two sections of the device. Fits of EDCs through KW reveal CBM energies of -0.08 eV and -0.03 eV in the vacuum exposed and graphene-encapsulated parts, respectively (see Fig. 2(c)). The EDC linewidth is nearly doubled in bare SL WS2 due to extrinsic broadening from the photocurrent. Considering a quadratic dispersion, we estimate the SL WS2 carrier concentration $n_{\mathrm{W}}$ from the fitted CBM. The values of $n_{\mathrm{W}}$ are stated in Figs. 2(a)-(b) and reveal a doping that is more than twice as large in vacuum exposed SL WS2 at $V_{\mathrm{G}}=8$ V than in the graphene-encapsulated part at $V_{\mathrm{G}}=12$ V.
@@
Figure 3: (a)-(b) $V_{\mathrm{G}}$-dependent sweeps of EDC at $\mathrm{K}_{\mathrm{W}}$ of SL WS2 measured on (a) vacuum exposed and (b) graphene-encapsulated sections. The $V_{\mathrm{G}}$ ranges are restricted to where intensity from the CBM emerges in the two sections. (c) Mean energies of VBM and CBM positions of SL WS2 for vacuum exposed (red markers) and graphene-encapsulated (blue markers) sections across the complete measured $V_{\mathrm{G}}$ range between -2 V and 12 V. Dashed green and orange boxes demarcate the $V_{\mathrm{G}}$- and energy-ranges presented in (a) and (b). Double-headed arrows demarcate band gaps $\mathrm{\Delta}$. Mean values and standard deviation of the gaps in the two sections of SL WS2 are stated in the band diagrams.
@@
By repeatedly ramping $V_{\mathrm{G}}$ up and down between -2 V to 12 V, we continuously and reversibly tune the band positions, which we extract from EDCs as in Figs. 3(a)-(b) in both sections of the device, as summarized in Fig. 3(c). Note that increasing energy broadening makes extraction of these band positions impossible above 10 V on vacuum exposed WS2. From the extracted VBM and CBM positions we obtain estimates of the direct band gap $\mathrm{\Delta}$ in the two sections of the device. Surprisingly, we observe that $\mathrm{\Delta}=(2.03\pm 0.02)$ eV in vacuum exposed WS2 while we find that $\mathrm{\Delta}=(2.19\pm 0.03)$ eV in graphene-encapsulated WS2 across the measured $V_{\mathrm{G}}$ range (see band diagrams in Fig. 3(c)). Subtle changes of $\mathrm{\Delta}$ with $V_{\mathrm{G}}$ occur, but they are within the experimental uncertainty [34]. Considering only the additional environmental screening from the graphene layer one would expect the opposite trend of what we observe: The overall Coulomb interaction should be suppressed in the graphene-encapsulated area yielding a reduced local quasiparticle gap [9, 11, 13, 37].
In a more detailed analysis, the vacuum exposed section of the device should be understood as a situation with a small environmental screening but large electron doping level, while the graphene-encapsulated section corresponds to a large environmental screening and small electron doping level. To disentangle these internal and external screening channels at finite charge doping by theoretical means, we performed ab initio model $G_{0}W_{0}$ calculations for WS2 under the influence of electron doping ($n_{\mathrm{W}}>0$) and environmental dielectric screening ($\varepsilon_{\mathrm{env}}>1)$ [34].
@@
In Fig. 4(a) we show an example of a $G_{0}W_{0}$ spectral function for a small electron doping ($n_{\mathrm{W}}\approx 1.82\times 10^{13}$ cm-2) and an intermediate environmental screening ($\varepsilon_{\mathrm{env}}=5$) around $\mathrm{K}_{\mathrm{W}}$. The quasiparticle VBM $E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$ occurs around $-2$ eV (spin-orbit coupling effects are neglected) and the CBM is occupied, such that $E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$ (red arrow in Fig. 4(a)) is below 0 eV. As a result of the finite doping, low-energetic (“acoustic”) plasmons are formed, which can dress the quasiparticles yielding plasmon satellite features shifted by a plasma energy $\omega_{\mathrm{pl}}(\mathrm{K}_{\mathrm{W}})$ (see purple arrow in Fig. 4(a)) [37, 38, 39]. In Fig. 4(b) we show the corresponding EDC taken at $\mathrm{K}_{\mathrm{W}}$ next to EDCs for different doping levels and a standard undoped DFT calculation for reference. In the following, we differentiate between the quasiparticle gap $\mathrm{\Delta}_{\mathrm{qp}}=E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})-E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$, and an “effective gap” that we define as $\mathrm{\Delta}_{\mathrm{e}}=E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})-\frac{1}{2}\omega_{\mathrm{pl}}(\mathrm{K}_{\mathrm{W}})-E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$, which are indicated by double-headed arrows in Fig. 4(b). In Figs. 4(c)-(d) we further show how $\mathrm{\Delta}_{\mathrm{qp}}$ and $\mathrm{\Delta}_{\mathrm{e}}$ are affected by $n_{\mathrm{W}}$ and $\varepsilon_{\mathrm{env}}$.
From Fig. 4(c) we see that $\mathrm{\Delta}_{\mathrm{qp}}$ decreases by only about 25 meV with $n_{\mathrm{W}}$ increasing from $1.82\times 10^{13}$ cm-2 to $6.14\times 10^{13}$ cm-2. This results from just minor changes to the density of states at the Fermi level, such that the Thomas-Fermi screening length and the final Coulomb interaction only mildly change. Increasing $\varepsilon_{\mathrm{env}}$ from $1$ to $9$ decreases $\mathrm{\Delta}_{\mathrm{qp}}$ by more than 60 meV. This trend is mainly controlled by the (effective) thickness of WS2 and thus by its distance to the dielectric environment. It is the ratio of the Thomas-Fermi screening length to this (effective) thickness that mostly controls the relative impact of $n_{\mathrm{W}}$ and $\varepsilon_{\mathrm{env}}$ to $\mathrm{\Delta}_{\mathrm{qp}}$ [25]: The smaller the density of states at the Fermi level (the smaller the Thomas-Fermi screening length) and the smaller the effective thickness of the 2D material, the larger the impact of $\varepsilon_{\mathrm{env}}$ to $\mathrm{\Delta}_{\mathrm{qp}}$.
@@
Our minimal many-body modeling reveals an intriguing interplay between external and internal screening channels (here enabled by the relatively low doping level), that is responsible for larger quasiparticle gaps in areas with large external screening but reduced doping levels and smaller quasiparticle gaps in areas with smaller external screening but larger doping levels. The ARPES observations are thus driven by modulated screening effects uncovering the relevance of spatially varying many-body effects in device architectures based on layered materials.
0.1 Acknowledgements
C.S. acknowledges Marie Sklodowska-Curie Postdoctoral Fellowship (project 101059528 MaPWave). The authors acknowledge funding from the Novo Nordisk Foundation (Project Grant NNF22OC0079960, and NFF23OC0085585), European Research Council (ERC) grant EXCITE (Grant. No. 101124619), VILLUM FONDEN under the Villum Investigator Program (Grant. No. 25931), the Danish Council for Independent Research (Grant Nos. DFF-9064-00057B, DFF-6108-00409, DFF-4258-00002B and 1026-00089B), and the Aarhus University Research Foundation. Y.V. and M.R. acknowledge support from the Dutch Research Council (NWO) via the “TOPCORE” consortium. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) , the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan. C.S. acknowledges Tony F. Heinz for helpful discussions.

### v2

# Quasiparticle gap renormalization driven by internal and external screening in a WS2 device

@@ Abstract
Figure 1: (a) Simplified schematic of microARPES experiment on a gated SL WS2 device. Gate and drain electrodes are labelled as G and D. (b) Optical image of the device. Note that gate and drain electrodes have been pre-patterned and are therefore located under graphite and graphene, respectively. The setup is equivalent to the simplified diagram in (a). (c) Energy- and momentum-integrated map of ARPES intensity from the device region. The red triangular outline and the blue square outline in (b)-(c) demarcate the SL WS2 and graphene flakes, respectively. (d) ARPES spectra of the graphene Dirac cone at the given gate voltages ($V_{G}$) obtained along the cut through $\mathrm{K}_{\mathrm{Gr}}$ marked on the BZs of graphene (black hexagon) and SL WS2 (red hexagon) in the inset. The measurement was performed in the spot marked by a blue star in (c). (e) Graphene carrier density $n_{\mathrm{Gr}}$ as a function of $V_{\mathrm{G}}$ obtained from ARPES spectra of the Dirac cone as shown in (d) [34]. Red full lines represent linear fits. The saturation level for electron-doping of the graphene is demarcated by a blue dashed line. (f)-(g) ARPES spectra of SL WS2 at the stated values of $V_{G}$ measured along $\mathrm{\Gamma}-\mathrm{K}_{\mathrm{W}}$. The ARPES cut direction is shown on the BZs in the inset. The energies of the VBM and CBM obtained from EDC fits are marked by red and blue ticks for spectra obtained in the two regions of the device demarcated by (f) red and (g) blue stars in (b)-(c) [34].
@@
Figure 2: (a)-(b) Close-up of ARPES spectra of the CBM region in (a) vacuum exposed and (b) graphene-encapsulated SL WS2 with estimates of electron density $n_{\mathrm{W}}$. (c) EDCs extracted along the dashed vertical line at $\mathrm{K}_{\mathrm{W}}$ in (a)-(b). Fits to a function composed of a single Lorentzian peak on a linear background multiplied by the Fermi-Dirac distribution are overlaid. The blue and red ticks demarcate the fitted peak positions. Double-headed arrows demarcate the full width at half maximum of the fitted Lorentzian peaks. Values are stated in units of eV and have error bars of ±0.02$\pm$ 0.02± 0.02 eV.
@@
Turning to the vacuum exposed section of SL WS2 in the device, we present the $E(k)$ band dispersion along $\mathrm{\Gamma}-\mathrm{K}_{\mathrm{W}}$ in Fig. 1(f). Note that the Fermi level is always referenced to the Fermi energy of the graphene top contact. We present spectra at $V_{\mathrm{G}}=0$ V and at the electron-doping saturation level of $V_{\mathrm{G}}=8$ V determined on graphene. At $V_{\mathrm{G}}=0$ V, we observe the valence band maximum (VBM) to be located at $\mathrm{K}_{\mathrm{W}}$ at an energy of -1.48 eV and with a spin-orbit splitting of 0.42 eV, consistent with the expected band structure of SL WS2 [36]. At $V_{\mathrm{G}}=8$ V, the VBM is shifted by 0.61 eV to lower energies, the CBM has shifted below $E_{\mathrm{F}}$ and a substantial energy broadening of the bands has occurred. We find that the linewidth of energy distribution curves (EDCs) through the VBM nearly doubles compared to the situation at $V_{\mathrm{G}}=0$ V [34]. Such broadening primarily results from photo-induced currents in the SL WS2 that leads to electrostatic potential variations across the 4 $\mu$m length scale of the light spot, which causes inhomogeneous doping and energy-shifting of the bands [27, 30, 31].
@@
We do not encounter this problem in the graphene-encapsulated section of the SL WS2 device, as we do not observe any energy broadening under the graphene [34]. In this case, the photoemitted charge in SL WS2 is efficiently replenished via the highly conductive graphene contact, which greatly reduces inhomogeneous potential variations across the light spot. ARPES spectra are presented in Fig. 1(g). The reduced intensity and higher background compared with the spectra of bare SL WS2 in Fig. 1(f) are caused by inelastic photoelectron scattering in the overlying graphene. At $V_{\mathrm{G}}=0$ V, the VBM is located at -1.40 eV and thus shifted further towards $E_{\mathrm{F}}$ than in the vacuum exposed SL WS2 section of the device. This shift is due to the enforced band alignment upon interfacing SL WS2 with graphene. Tuning to $V_{\mathrm{G}}=12$ V leads to a downwards VBM shift of 0.80 eV and a faint CBM state at $E_{\mathrm{F}}$. We do not observe additional energy and momentum broadening of the SL WS2 bands in this part of the device when we increase $V_{\mathrm{G}}$ to the maximum possible value of 12 V before gate leakage through hBN sets in [34].
@@
Detailed ARPES spectra of the SL WS2 CBM region are presented in Fig. 2 for the two sections of the device. Fits of EDCs through KW reveal CBM energies of -0.08 eV and -0.03 eV in the vacuum exposed and graphene-encapsulated parts, respectively (see Fig. 2(c)) [34, 37, 38]. The EDC linewidth is nearly doubled in bare SL WS2 due to extrinsic broadening from the photocurrent. Considering a quadratic dispersion, we estimate the SL WS2 carrier concentration $n_{\mathrm{W}}$ from the fitted CBM. The values of $n_{\mathrm{W}}$ are stated in Figs. 2(a)-(b) and reveal a doping that is more than twice as large in vacuum exposed SL WS2 at $V_{\mathrm{G}}=8$ V than in the graphene-encapsulated part at $V_{\mathrm{G}}=12$ V.
@@
Figure 3: (a)-(b) $V_{\mathrm{G}}$-dependent sweeps of EDC at $\mathrm{K}_{\mathrm{W}}$ of SL WS2 measured on (a) vacuum exposed and (b) graphene-encapsulated sections. The $V_{\mathrm{G}}$ ranges are restricted to where intensity from the CBM emerges in the two sections. (c) Mean energies of VBM and CBM positions of SL WS2 for vacuum exposed (red markers) and graphene-encapsulated (blue markers) sections across a wide $V_{\mathrm{G}}$ range between -2 V and 12 V. Note that measurements were performed down to $V_{\mathrm{G}}=-6$ V, which are not shown since the bands merely exhibit a monotonic shift with gate voltage in this range. Dashed green and orange boxes demarcate the $V_{\mathrm{G}}$- and energy-ranges presented in (a) and (b). Double-headed arrows demarcate band gaps $\mathrm{\Delta}$. Mean values and standard deviation of the gaps in the two sections of SL WS2 are stated in the band diagrams.
@@
By repeatedly ramping $V_{\mathrm{G}}$ up and down between negative and positive voltage values, we continuously and reversibly tune the band positions, which we extract from EDCs as in Figs. 3(a)-(b) in both sections of the device, as summarized in Fig. 3(c). Note that increasing energy broadening makes extraction of these band positions impossible above 10 V on vacuum exposed WS2. From the extracted VBM and CBM positions we obtain estimates of the direct band gap $\mathrm{\Delta}$ in the two sections of the device. Surprisingly, we observe that $\mathrm{\Delta}=(2.03\pm 0.02)$ eV in vacuum exposed WS2 while we find that $\mathrm{\Delta}=(2.19\pm 0.03)$ eV in graphene-encapsulated WS2 across the measured $V_{\mathrm{G}}$ range (see band diagrams in Fig. 3(c)). Subtle changes of $\mathrm{\Delta}$ with $V_{\mathrm{G}}$ occur, but they are within the experimental uncertainty [34]. Considering only the additional environmental screening from the graphene layer one would expect the opposite trend of what we observe: The overall Coulomb interaction should be suppressed in the graphene-encapsulated area yielding a reduced local quasiparticle gap [9, 11, 13, 39].
In a more detailed analysis, the vacuum exposed section of the device should be understood as a situation with a small environmental screening but large electron doping level, while the graphene-encapsulated section corresponds to a large environmental screening and small electron doping level. To disentangle these internal and external screening channels at finite charge doping by theoretical means, we performed ab initio model $G_{0}W_{0}$ calculations for WS2 under the influence of electron doping ($n_{\mathrm{W}}>0$) and environmental dielectric screening ($\varepsilon_{\mathrm{env}}>1)$ [34, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
@@
In Fig. 4(a) we show an example of a $G_{0}W_{0}$ spectral function for a small electron doping ($n_{\mathrm{W}}\approx 1.82\times 10^{13}$ cm-2) and an intermediate environmental screening ($\varepsilon_{\mathrm{env}}=5$) around $\mathrm{K}_{\mathrm{W}}$. The quasiparticle VBM $E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$ occurs around 2-2- 2 eV (spin-orbit coupling effects are neglected) and the CBM is occupied, such that $E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$ (red arrow in Fig. 4(a)) is below 0 eV. As a result of the finite doping, low-energetic (“acoustic”) plasmons are formed, which can dress the quasiparticles yielding plasmon satellite features shifted by a plasma energy $\omega_{\mathrm{pl}}(\mathrm{K}_{\mathrm{W}})$ (see purple arrow in Fig. 4(a)) [39, 54, 55]. In Fig. 4(b) we show the corresponding EDC taken at $\mathrm{K}_{\mathrm{W}}$ next to EDCs for different doping levels and a standard undoped DFT calculation for reference. In the following, we differentiate between the quasiparticle gap $\mathrm{\Delta}_{\mathrm{qp}}=E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})-E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$, and an “effective gap” that we define as $\mathrm{\Delta}_{\mathrm{e}}=E^{(\mathrm{c})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})-\frac{1}{2}\omega_{\mathrm{pl}}(\mathrm{K}_{\mathrm{W}})-E^{(\mathrm{v})}_{\mathrm{qp}}(\mathrm{K}_{\mathrm{W}})$, which are indicated by double-headed arrows in Fig. 4(b). In Figs. 4(c)-(d) we further show how $\mathrm{\Delta}_{\mathrm{qp}}$ and $\mathrm{\Delta}_{\mathrm{e}}$ are affected by $n_{\mathrm{W}}$ and $\varepsilon_{\mathrm{env}}$.
From Fig. 4(c) we see that $\mathrm{\Delta}_{\mathrm{qp}}$ decreases by only about 25 meV with $n_{\mathrm{W}}$ increasing from $1.82\times 10^{13}$ cm-2 to $6.14\times 10^{13}$ cm-2. This results from just minor changes to the density of states at the Fermi level, such that the Thomas-Fermi screening length and the final Coulomb interaction only mildly change. Increasing $\varepsilon_{\mathrm{env}}$ from 111 to 999 decreases $\mathrm{\Delta}_{\mathrm{qp}}$ by more than 60 meV. This trend is mainly controlled by the (effective) thickness of WS2 and thus by its distance to the dielectric environment. It is the ratio of the Thomas-Fermi screening length to this (effective) thickness that mostly controls the relative impact of $n_{\mathrm{W}}$ and $\varepsilon_{\mathrm{env}}$ to $\mathrm{\Delta}_{\mathrm{qp}}$ [25]: The smaller the density of states at the Fermi level (the smaller the Thomas-Fermi screening length) and the smaller the effective thickness of the 2D material, the larger the impact of $\varepsilon_{\mathrm{env}}$ to $\mathrm{\Delta}_{\mathrm{qp}}$.
@@
Our minimal many-body modeling reveals an intriguing interplay between external and internal screening channels (here enabled by the relatively low doping level), that is responsible for larger quasiparticle gaps in areas with large external screening but reduced doping levels and smaller quasiparticle gaps in areas with smaller external screening but larger doping levels. The ARPES observations are thus driven by modulated screening effects uncovering the relevance of spatially varying many-body effects in device architectures based on layered materials. [56]
1 Acknowledgements
C.S. acknowledges Marie Sklodowska-Curie Postdoctoral Fellowship (project 101059528 MaPWave). The work was funded/co-funded by the European Union (ERC grant EXCITE with project number 101124619). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors acknowledge funding from the Novo Nordisk Foundation (Project Grant NNF22OC0079960, and NFF23OC0085585), VILLUM FONDEN under the Villum Investigator Program (Grant. No. 25931), the Danish Council for Independent Research (Grant Nos. DFF-9064-00057B, DFF-6108-00409, DFF-4258-00002B and 1026-00089B), and the Aarhus University Research Foundation. Y.V. and M.R. acknowledge support from the Dutch Research Council (NWO) via the “TOPCORE” consortium. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) , the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan. C.S. acknowledges Tony F. Heinz for helpful discussions.